Pharmacology 5. Order: Ampicillin-Sulbactam 250 mg/IV based on ampicillin content. Stock: Ampicillin-Sulbactam 500:250 mg/vial. Available:
Answer #1

Concentration = 45 mg/ml

1500 mg in = 1/45 * 1500 = 33.3 ml

So 33 ml diluent should be added.

1500 mg in 33 ml

250 mg in 33/1500 * 250 = 5.5 ml.